Ask
your own question, for FREE!
Mathematics
13 Online
Help! Will Give Medal Nine door prizes are to be awarded at a club meeting. There are 100 tickets in the box, and one of them is yours. A ticket is drawn, the prize is awarded, and the ticket is placed back in the box. This continues until all nine prizes are awarded. (a) Find the probability your ticket is drawn once. (b) Find the probability your ticket is drawn twice.
Still Need Help?
Join the QuestionCove community and study together with friends!
I'm going to have to guess here... 1/100 = .01 .01/9 .02/9 would be the second answer by that logic.
a) there are (9C1) ways your ticket can be drawn. So P = (9C1) (1/100) (99/100)^8 b) same argument P = (9C2) (1/100)^2 (99/100)^7
Thank You
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
7 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal